# Francisco Calvillo LPSM, Sorbonne Université, 4 Place Jussieu, 75005 Paris, France Luc Devroye

### v1

@@ 2 The one-choice model
Assume $\mathbb{E}Z<\infty$ and $q_{1}>0~{}.$ Then there exists a positive random variable $N$ with $\mathbb{P}\left(N<\infty\right)=1$ such that, with probability one,
@@
Note that the actual statement of the theorem in [6] is slightly weaker since it only asserts that there is a random variable $N$, finite almost surely, such that for all $n\geq N$, the vertex $n$ belongs to the same tree as the vertex $N$. This does not necessarily imply that $N\in T_{n}^{Z}$. Nevertheless, one can add this property of $N$ without changing the proof provided in [6].
Define $Z^{\prime}_{n}=\min(Z_{n},W_{n})$ and $Z^{\prime}=\min(Z,W)$ and note that $S_{n}=T_{n}^{Z^{\prime}}$. By Theorem 8 we obtain the following result for moderate and light tails.
@@
If $\mathbb{E}\min(Z,W)<\infty$ and $q_{1}>0$, then there exists a positive random variable $N$ with $\mathbb{P}\left(N<\infty\right)=1$ such that, with probability one,
$N\in S_{n}\quad\textit{for all }n\geq N$
@@
If $\mathbb{E}Z=\infty$ and $q_{i}>0$ for all $i\geq 1$, then
@@ 3 Rightmost and leftmost leaves
As stated in Theorem 4, the leftmost leaf $M_{n}$ reached by vertex $n$ diverges to $\infty$- ∞ whenever $\mathbb{E}Z=\infty$. This can be seen as a consequence of the behavior of the one-choice model, as one only needs to check the divergence to negative infinity of the leaf $\min T_{n}^{Z}$ reached by vertex $n$ by following only the $Z$-edges.
Proof.  [Proof of Theorem 4] From Theorem 10 we know that with probability one, for all $k\leq 0$, there are at most a finite number of integers $n\geq 1$ such that $k\in T_{n}^{Z}$. Thus, the sequence $(\min T_{n}^{Z})_{n\in\mathbb{N}}$, taking values in $\{0,-1,-2,\dots\}$, cannot take the same value an infinite number of times, so it goes to $\infty$- ∞ as $n\to\infty$ with probability one. This concludes the proof, since $M_{n}\leq\min T_{n}^{Z}$ for all $n\geq 1$.
@@
$\sum_{n\geq 1}\mathbb{P}\left(\min(Z_{n},W_{n})\geq 2n\right)=\sum_{n\geq 1}p_{2n}^{2}=\infty~{}.$
@@ 4.1 Moderate tails
Consider the number $L_{n}$ of leaves reached by vertex $n$, when the distribution of $Z$ is such that $\mathbb{E}Z=\infty$ and $\mathbb{E}\min(Z,W)<\infty$. We first show weak divergence, implying weak consistency in configuration (iii).
@@
Assume that the distribution of $Z$ is not supported by any proper additive subgroup of the integers. If $\mathbb{E}Z=\infty$ and $\mathbb{E}\min(Z,W)<\infty$, then
$L_{n}\to\infty~{}\textit{ in probability as }n\to\infty.$
@@
Define
$J_{n}=\sum_{m=1}^{n}\mathbbm{1}_{m\in S_{n}}\mathbbm{1}_{\max(Z_{m},W_{m})\geq m}~{}.$
Then $\mathbb{E}J_{n}\to\infty$ as $n\to\infty$.
Proof.   $J_{n}$ counts the number of vertices in $S_{n}$ with a positive index that are connected to a leaf through its longest outgoing edge. Note that for $n\geq m\geq 1$, the events $\{m\in S_{n}\}$ and $\{\max(Z_{m},W_{m})\geq m\}$ are independent. It follows that
$\mathbb{E}J_{n}=\sum_{m=1}^{n}\mathbb{P}(m\in S_{n})\mathbb{P}(\max(Z_{m},W_{m})\geq m)=\sum_{m=1}^{n}v_{n-m}\mathbb{P}(\max(Z_{m},W_{m})\geq m)$
where $v_{m}:=\mathbb{P}(0\in S_{m})$ for all $m\geq 1$, and $v_{0}=1$.
Let $f(z)=\sum_{n\geq 0}v_{n}z^{n}$, $\tilde{q}_{m}=\mathbb{P}(\min(Z,W)=m)$ and $g(z)=\sum_{n\geq 1}\tilde{q}_{n}z^{n}$. In view of
$v_{n}=\sum_{m=0}^{n-1}\tilde{q}_{n-m}v_{m},$
we have
$f(z)=1+\sum_{n=1}^{\infty}z^{n}\sum_{m=0}^{n}\tilde{q}_{n-m}v_{m}=1+g(z)f(z)=\frac{1}{1-g(z)}.$
Since $\min(Z,W)$ is not supported by any proper additive subgroup of the integers, the Erdős-Feller-Pollard theorem [9] implies that
$\lim_{n\to\infty}v_{n}=\frac{1}{\mathbb{E}\min(Z,W)}~{}.$ (4.1)
@@
$\mathbb{E}J_{n}\geq c\sum_{m=1}^{n}p_{m}\to\infty\quad\text{as }n\to\infty~{}.$ (4.2)
@@
Note that for $n\geq 1$,
$(\mathbb{E}J_{n})^{2}$ $=2\sum_{1\leq k<m\leq n}p_{m}^{*}p_{k}^{*}v_{n-m}v_{n-k}+\sum_{1\leq m\leq n}(p_{m}^{*})^{2}v_{n-m}^{2}$
$=2\sum_{1\leq k<m\leq n}p_{m}^{*}p_{k}^{*}v_{n-m}v_{n-k}+O(1)$ (4.4)
since $\sum_{1\leq m\leq n}(p_{m}^{*})^{2}v_{n-m}^{2}$ is bounded by $4\mathbb{E}\min(Z,W)<\infty$.
@@
$\mathbb{E}\left[J_{n}^{2}\right]$ $=2\sum_{1\leq k<m\leq n}\mathbb{P}\left(m\in S_{n},k\in S_{n},X_{m}\geq m,X_{k}\geq k\right)+\mathbb{E}J_{n}$
$=2\sum_{1\leq k<m\leq n}p_{k}^{*}\mathbb{P}\left(m\in S_{n},k\in S_{n},X_{m}\geq m\right)+\mathbb{E}J_{n}$
$=2\sum_{1\leq k<m\leq n}p_{k}^{*}\mathbb{P}\mathopen{}\left[k\in S_{n},X_{m}\geq m\middle|m\in S_{n}\right]\mathbb{P}\left(m\in S_{n}\right)+\mathbb{E}J_{n}$
If $m\in S_{n}$, the event $\{k\in S_{n},X_{m}\geq m\}$ can only happen if one of the two random variables $Z_{m}$ and $W_{m}$ is greater or equal to $m$ and the other one is smaller than $m-k$, for otherwise $k$ could not belong to $S_{n}$ when $1\leq k<m\leq n$. Thus, using the independence of $Z_{m}$ with respect to $W_{m}$, $\{k\in S_{n}\}$ and $\{m\in S_{n}\}$, and using the fact that $p_{m}^{*}+p_{m}^{2}=2p_{m}$,
@@
$=2\sum_{1\leq k<m\leq n}p_{k}^{*}\times 2\mathbb{P}\mathopen{}\left[Z_{m}\geq m,W_{m}\leq m-k,k\in S_{n}\middle|m\in S_{n}\right]\mathbb{P}\left(m\in S_{n}\right)+\mathbb{E}J_{n}$
$=2\sum_{1\leq k<m\leq n}p_{k}^{*}(p_{m}^{*}+p_{m}^{2})v_{n-m}\mathbb{P}\mathopen{}\left[W_{m}\leq m-k,k\in S_{n}\middle|m\in S_{n}\right]+\mathbb{E}J_{n}$
$=4\sum_{1\leq k<m\leq n}p_{k}^{*}(p_{m}^{*}+p_{m}^{2})v_{n-m}\sum_{i=1}^{k-m}\mathbb{P}\left(k\in S_{m-i}\right)\mathbb{P}\left(W_{m}=i\right)+\mathbb{E}J_{n}$
$=2\sum_{1\leq k<m\leq n}p_{k}^{*}(p_{m}^{*}+p_{m}^{2})v_{n-m}\sum_{i=1}^{m-k}v_{m-k-i}q_{i}+\mathbb{E}J_{n}~{}.$ (4.6)
@@
$\sum_{1\leq k<m\leq n}p_{k}^{*}p_{m}^{2}v_{n-m}\sum_{i=1}^{m-k}q_{i}v_{m-k-i}$ $\leq\sum_{1\leq k<m\leq n}p_{k}^{*}p_{m}^{2}v_{n-m}$
$\leq\sum_{1\leq k<m\leq n}p_{k}^{*}p_{m}^{2}$
$\leq\left(\sum_{m=1}^{\infty}p_{m}^{2}\right)\left(\sum_{k=1}^{n}p_{k}^{*}\right)$
@@
$\mathbb{E}[J_{n}^{2}]=2\sum_{1\leq k<m\leq n}p_{k}^{*}p_{m}^{*}v_{n-m}\sum_{i=1}^{m-k}v_{m-k-i}q_{i}+O\left(\mathbb{E}J_{n}\right)~{}.$ (4.7)
@@
$\sum_{i=1}^{m}q_{i}v_{m-i}$ $\leq\sum_{i=1}^{m-x}q_{i}(\lambda+\epsilon)+\sum_{m-x<i\leq m}q_{i}$
@@
$\sum_{i=1}^{m}q_{i}v_{m-i}\leq\lambda+2\epsilon~{}.$
@@
$\sum_{\begin{subarray}{c}1\leq k<m\leq n\\ \textit{s.t. }m-k<y\end{subarray}}p_{m}^{*}v_{n-m}\underbrace{p_{k}^{*}\sum_{i=1}^{m-k}v_{m-k-i}q_{i}}_{\leq 1}$ $\leq\sum_{1\leq m\leq n}p_{m}^{*}v_{n-m}\sum_{k=m-y}^{m}1$
@@
$\sum_{\begin{subarray}{c}1\leq k<m\leq n\\ \textit{s.t. }n-m<y\end{subarray}}p_{k}^{*}\underbrace{p_{m}^{*}v_{n-m}\sum_{i=1}^{m-k}v_{m-k-i}q_{i}}_{\leq 1}$ $\leq\sum_{n-y\leq m\leq n}1\sum_{k=1}^{n}p_{k}^{*}$
$\leq(y+1)\sum_{k=1}^{n}p_{k}^{*}=O\left(\mathbb{E}J_{n}\right)~{}.$
@@
$\mathbb{E}[J_{n}^{2}]=\sum_{m=1}^{n-y}\sum_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}v_{n-k}\sum_{i=1}^{m-k}q_{i}v_{m-k-i}+O\left(\mathbb{E}J_{n}\right)~{}.$ (4.8)
@@
$(\mathbb{E}J_{n})^{2}=\sum_{s=1}^{n-y}\sum_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}v_{n-m}v_{n-k}+O\left(\mathbb{E}J_{n}\right)~{}.$ (4.9)
@@
$\sum_{m=1}^{n-y}\sum_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}\underbrace{v_{n-m}\sum_{i=1}^{m-k}q_{i}v_{m-k-i}}_{\leq(\lambda+2\epsilon)^{2}}\leq(\lambda+2\epsilon)^{2}\sum_{m=1}^{n-y}\sum_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}$ (4.10)
@@
$\sum_{m=1}^{n-y}\sum_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}\underbrace{v_{n-m}v_{n-k}}_{\geq(\lambda-\epsilon)^{2}}\geq(\lambda-\epsilon)^{2}\sum_{m=1}^{n-y}\sum_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}~{}.$ (4.11)
@@
Proposition 11 follows from the fact that $J_{n}\to\infty$ in probability. In order to see this, we only need to check that the fact that $J_{n}\to\infty$ in probability implies that the number of leaves connected to $T_{n}$ goes to infinity as well.
@@
$\leq\mathbb{P}\left(\left[\max_{k\in\mathcal{L}_{n}^{*}}\sum_{m\in S_{n}}\mathbbm{1}_{\max(Z_{m},W_{m})=m-k}<\frac{\mathbb{E}J_{n}}{2x}\right]\cap\left[\left|\mathcal{L}_{n}^{*}\right|<x\right]\right)$
@@
$J_{n}=\sum_{k\in\mathcal{L}_{n}^{*}}\sum_{m\in S_{n}}\mathbbm{1}_{\max(Z_{m},W_{m})=m-k}~{},$
@@
$\mathbb{P}\left(\left[\max_{k\in\mathcal{L}_{n}^{*}}\sum_{m\in S_{n}}\mathbbm{1}_{\max(Z_{m},W_{m})=m-k}<\frac{\mathbb{E}J_{n}}{2x}\right]\cap\left[\left|\mathcal{L}_{n}^{*}\right|<x\right]\right)$
@@
$\mathbb{P}(D_{n,k}>y)$ $=\mathbb{P}\left(\sum_{m=1}^{n}\mathbbm{1}_{\max(Z_{m},W_{m})=m-k}>y\right)$
@@
$\sum_{m=1}^{\infty}(p_{m}^{*})^{y_{n}}\leq\sum_{m=1}^{\infty}(p_{m}^{*})^{2}<\infty\text{ (since }\mathbb{E}\min(Z,W)<\infty\text{ )}~{}.$
@@
$\mathbb{P}\left(\cup_{k\leq 0}\left[D_{n,k}\geq\frac{\mathbb{E}J_{n}}{2x}\right]\right)$ $\leq\sum_{k\leq 0}\left(\frac{e}{y_{n}}\right)^{y_{n}}(p_{1-k}^{*})^{y_{n}}$
$\leq\left(\frac{e}{y_{n}}\right)^{y_{n}}\sum_{k=1}^{\infty}(p_{k}^{*})^{2}\to 0~{}.$ (4.15)
@@
$\mathbb{P}\left(Y_{n}\leq k\right)\leq\sum_{m\geq n}\mathbb{P}\left(\min(Z_{m},W_{m})>m-k\right).$
Since $\mathbb{E}\min(Z,W)<\infty$, the sum above is finite and goes to 00 as $n$ goes to infinity. Hence, $Y_{n}\to\infty$ in probability.
Let us fix $n\geq 0$ and apply Corollary 9 to the model obtained by shifting by $n$ the set of integers so that leaves are indexed by the set $\{k:k\leq n\}$ and the edges are given by the sequences of random variables $(Z_{k})_{k\geq n+1}$ and $(W_{k})_{k\geq n+1}$. This gives us a random variable $N_{n}>n$, finite almost surely, such that $N_{n}\in S_{k}$ for all $k\geq N_{n}$. Let $I_{n}\leq n\}$ be the root of this tree and note that

### v2

@@ 2 The one-choice model
Let $Z$ and $(Z_{n})_{n\geq 1}$ be i.i.d. random variables. Assume $\mathbb{E}Z<\infty$ and $q_{1}>0~.$ Then there exists a positive random variable $N$ with $\mathbb{P}\left(N<\infty\right)=1$ such that, with probability one,
@@
Note that the actual statement of the theorem in [6] is slightly weaker, as it only asserts the existence of a random variable $N$, finite almost surely, such that for all $n\geq N$, the vertex $n$ has the same colour as vertex $N$. This does not necessarily imply that $N\in T_{n}^{Z}$. Nevertheless, one can add this property of $N$ without changing the proof provided in [6].
Define $Z^{\prime}_{n}=\min(Z_{n},W_{n})$ and $Z^{\prime}=\min(Z,W)$ and note that $S_{n}=T_{n}^{Z^{\prime}}$. By Theorem 8, we obtain the following result for moderate and light tails.
@@
Let $Z$, $W$, $(Z_{n})_{n\geq 1}$ and $(W_{n})_{n\geq 1}$ be i.i.d. random variables. If $\mathbb{E}\min(Z,W)<\infty$ and $q_{1}>0$, then there exists a positive random variable $N$ with $\mathbb{P}\left(N<\infty\right)=1$ such that, with probability one,
$N\in S_{n}\quad\textit{for all }n\geq N~.$
@@
Let $Z$ and $(Z_{n})_{n\geq 1}$ be i.i.d. random variables. If $\mathbb{E}Z=\infty$ and $q_{i}>0$ for all $i\geq 1$, then
@@ 3 Rightmost and leftmost leaves
As stated in Theorem 4, the leftmost leaf $L_{n}$ reached by vertex $n$ diverges to -$\infty$ whenever $\mathbb{E}Z=\infty$. This can be seen as a consequence of the behavior of the one-choice model, as one only needs to check the divergence to negative infinity of the leaf $\min T_{n}^{Z}$ reached by vertex $n$ by following only the $Z$-edges.
Proof. [Proof of Theorem 4] From Theorem 10 we know that with probability one, for all $k\leq 0$, there are at most a finite number of integers $n\geq 1$ such that $k\in T_{n}^{Z}$. Thus, the sequence $(\min T_{n}^{Z})_{n\in\mathbb{N}}$, taking values in $\{0,-1,-2,\dots\}$, cannot take the same value an infinite number of times, so it goes to -$\infty$ as $n\to\infty$ with probability one. This concludes the proof, since $L_{n}\leq\min T_{n}^{Z}$ for all $n\geq 1$.
@@
$\sumop\displaylimits_{n\geq 1}\mathbb{P}\left(\min(Z_{n},W_{n})\geq 2n\right)=\sumop\displaylimits_{n\geq 1}p_{2n}^{2}=\infty~.$
@@ 4.1 Moderate tails
Consider the number n of leaves reached by vertex $n$, when the distribution of $Z$ is such that $\mathbb{E}Z=\infty$ and $\mathbb{E}\min(Z,W)<\infty$. We first show divergence in probability, implying weak consistency in configuration (iii).
@@
Assume that the distribution of $Z$ is aperiodic. If $\mathbb{E}Z=\infty$ and $\mathbb{E}\min(Z,W)<\infty$, then
[…]
The proof uses the balance that arises in the moderate-tail regime. The distribution of delays is sufficiently heavy to ensure that each non-leaf vertex in $T_{n}$ has a non-negligible chance of reaching a leaf, but light enough to guarantee that $T_{n}$ contains a substantial number of non-leaf vertices. Consequently, the overall number of opportunities to generate new leaves grows significantly with $n$. This balance is illustrated in Figure 4: only in the moderate tail regime ($\alpha=0.6$ in the figure) do we simultaneously observe a sizable number of blue vertices (non-leaves) and a large population of red vertices (leaves).
To formalize this heuristic, we focus on the chain $S_{n}$, obtained by repeatedly following the shortest outgoing edge starting from vertex $n$. In the moderate-tailed regime, this chain typically contains many vertices, and each of these vertices has an additional opportunity to connect to a leaf through its longest outgoing edge. We therefore define
$J_{n}=\sumop\displaylimits_{m=1}^{n}\mathbbm{1}_{m\in S_{n}}\,\mathbbm{1}_{\max(Z_{m},W_{m})\geq m},$
so that $J_{n}$ counts the number of vertices in $S_{n}$ that are connected to a leaf via their longest outgoing edge. Controlling the growth of $J_{n}$ makes it possible to turn the heuristic balance into a rigorous statement.
@@
The expected number of vertices in the chain $S_{n}$ that connect to a leaf through their longest outgoing edge grows to infinity with $n$:
$\mathbb{E}[J_{n}]\;\to\;\infty\quad\text{as }n\to\infty~.$
Proof. Note that for $n\geq m\geq 1$, the events $\{m\in S_{n}\}$ and $\{\max(Z_{m},W_{m})\geq m\}$ are independent. Define $v_{m}:=\mathbb{P}(0\in S_{m})$ for all $m\geq 1$ and let $v_{0}=1$, so that
$\mathbb{E}J_{n}=\sumop\displaylimits_{m=1}^{n}\mathbb{P}(m\in S_{n})\mathbb{P}(\max(Z_{m},W_{m})\geq m)=\sumop\displaylimits_{m=1}^{n}v_{n-m}\mathbb{P}(\max(Z_{m},W_{m})\geq m)~.$
The sequence $(v_{n})_{n\geq 0}$ can be better understood by considering the arithmetic renewal process $M_{n}$ with holding times given by the i.i.d. random variables $(\min(Z_{i},W_{i}))_{i\geq 1}$, and initialized with $M_{0}=0$. In other words,
$M_{n}=\sumop\displaylimits_{i=1}^{n}\min(Z_{i},W_{i})~.$
By definition, $v_{n}$ is the probability that $n$ can be expressed as a sum of independent random variables $(\min(Z_{i},W_{i}))_{i\geq 0}$. In terms of the renewal process, this means that
$v_{n}=\mathbb{P}(\exists k\geq 0:M_{k}=n)~.$
Hence, the sequence $(v_{n})_{n\geq 0}$ corresponds to the renewal measure of the aperiodic renewal process $(M_{n})_{n\geq 0}$. By the Erdős–Feller–Pollard renewal theorem [9], it follows that
$\lim_{n\to\infty}v_{n}=\frac{1}{\mathbb{E}\min(Z,W)}>0~.$ (4.1)
@@
$\mathbb{E}J_{n}\geq c\sumop\displaylimits_{m=1}^{n}p_{m}\to\infty\quad\text{as }n\to\infty~.$ (4.2)
@@
Recall that $v_{n}=\mathbb{P}(0\in S_{n})$, as introduced earlier. Then, for $n\geq 1$,
$(\mathbb{E}J_{n})^{2}$ $=2\sumop\displaylimits_{1\leq k<m\leq n}p_{m}^{*}p_{k}^{*}v_{n-m}v_{n-k}+\sumop\displaylimits_{1\leq m\leq n}(p_{m}^{*})^{2}v_{n-m}^{2}$
$=2\sumop\displaylimits_{1\leq k<m\leq n}p_{m}^{*}p_{k}^{*}v_{n-m}v_{n-k}+O(1)$ (4.4)
since $\sumop\displaylimits_{1\leq m\leq n}(p_{m}^{*})^{2}v_{n-m}^{2}$ is bounded by $4\mathbb{E}\min(Z,W)<\infty$.
@@
$\mathbb{E}\left[J_{n}^{2}\right]$ $=2\sumop\displaylimits_{1\leq k<m\leq n}\mathbb{P}\left(m\in S_{n},k\in S_{n},X_{m}\geq m,X_{k}\geq k\right)+\mathbb{E}J_{n}$
$=2\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}\mathbb{P}\left(m\in S_{n},k\in S_{n},X_{m}\geq m\right)+\mathbb{E}J_{n}$
$=2\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}\mathbb{P}\mathopen{}\left[k\in S_{n},X_{m}\geq m\middle|m\in S_{n}\right]\mathbb{P}\left(m\in S_{n}\right)+\mathbb{E}J_{n}$
If $m\in S_{n}$, the event $\{k\in S_{n},X_{m}\geq m\}$ can only happen if one of the two random variables $Z_{m}$ and $W_{m}$ is greater or equal to $m$ and the other one is smaller than mkm-k, for otherwise $k$ could not belong to $S_{n}$ when $1\leq k<m\leq n$. Thus, using the independence of $Z_{m}$ with respect to $W_{m}$, $\{k\in S_{n}\}$ and $\{m\in S_{n}\}$, and using the fact that $p_{m}^{*}+p_{m}^{2}=2p_{m}$,
@@
$=2\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}\times 2\mathbb{P}\mathopen{}\left[Z_{m}\geq m,W_{m}\leq m-k,k\in S_{n}\middle|m\in S_{n}\right]\mathbb{P}\left(m\in S_{n}\right)+\mathbb{E}J_{n}$
$=2\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}(p_{m}^{*}+p_{m}^{2})v_{n-m}\mathbb{P}\mathopen{}\left[W_{m}\leq m-k,k\in S_{n}\middle|m\in S_{n}\right]+\mathbb{E}J_{n}$
$=4\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}(p_{m}^{*}+p_{m}^{2})v_{n-m}\sumop\displaylimits_{i=1}^{k-m}\mathbb{P}\left(k\in S_{m-i}\right)\mathbb{P}\left(W_{m}=i\right)+\mathbb{E}J_{n}$
$=2\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}(p_{m}^{*}+p_{m}^{2})v_{n-m}\sumop\displaylimits_{i=1}^{m-k}v_{m-k-i}q_{i}+\mathbb{E}J_{n}~.$ (4.6)
@@
$\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}p_{m}^{2}v_{n-m}\sumop\displaylimits_{i=1}^{m-k}q_{i}v_{m-k-i}$ $\leq\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}p_{m}^{2}v_{n-m}$
$\leq\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}p_{m}^{2}$
$\leq\left(\sumop\displaylimits_{m=1}^{\infty}p_{m}^{2}\right)\left(\sumop\displaylimits_{k=1}^{n}p_{k}^{*}\right)$
@@
$\mathbb{E}[J_{n}^{2}]=2\sumop\displaylimits_{1\leq k<m\leq n}p_{k}^{*}p_{m}^{*}v_{n-m}\sumop\displaylimits_{i=1}^{m-k}v_{m-k-i}q_{i}+O\left(\mathbb{E}J_{n}\right)~.$ (4.7)
@@
$\sumop\displaylimits_{i=1}^{m}q_{i}v_{m-i}$ $\leq\sumop\displaylimits_{i=1}^{m-x}q_{i}(\lambda+\epsilon)+\sumop\displaylimits_{m-x<i\leq m}q_{i}$
@@
$\sumop\displaylimits_{i=1}^{m}q_{i}v_{m-i}\leq\lambda+2\epsilon~.$
@@
$\sumop\displaylimits_{\begin{subarray}{c}1\leq k<m\leq n\\ \textit{s.t. }m-k<y\end{subarray}}p_{m}^{*}v_{n-m}\underbrace{p_{k}^{*}\sumop\displaylimits_{i=1}^{m-k}v_{m-k-i}q_{i}}_{\leq 1}$ $\leq\sumop\displaylimits_{1\leq m\leq n}p_{m}^{*}v_{n-m}\sumop\displaylimits_{k=m-y}^{m}1$
@@
$\sumop\displaylimits_{\begin{subarray}{c}1\leq k<m\leq n\\ \textit{s.t. }n-m<y\end{subarray}}p_{k}^{*}\underbrace{p_{m}^{*}v_{n-m}\sumop\displaylimits_{i=1}^{m-k}v_{m-k-i}q_{i}}_{\leq 1}$ $\leq\sumop\displaylimits_{n-y\leq m\leq n}1\sumop\displaylimits_{k=1}^{n}p_{k}^{*}$
$\leq(y+1)\sumop\displaylimits_{k=1}^{n}p_{k}^{*}=O\left(\mathbb{E}J_{n}\right)~.$
@@
$\mathbb{E}[J_{n}^{2}]=\sumop\displaylimits_{m=1}^{n-y}\sumop\displaylimits_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}v_{n-k}\sumop\displaylimits_{i=1}^{m-k}q_{i}v_{m-k-i}+O\left(\mathbb{E}J_{n}\right)~.$ (4.8)
@@
$(\mathbb{E}J_{n})^{2}=\sumop\displaylimits_{s=1}^{n-y}\sumop\displaylimits_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}v_{n-m}v_{n-k}+O\left(\mathbb{E}J_{n}\right)~.$ (4.9)
@@
$\sumop\displaylimits_{m=1}^{n-y}\sumop\displaylimits_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}\underbrace{v_{n-m}\sumop\displaylimits_{i=1}^{m-k}q_{i}v_{m-k-i}}_{\leq(\lambda+2\epsilon)^{2}}\leq(\lambda+2\epsilon)^{2}\sumop\displaylimits_{m=1}^{n-y}\sumop\displaylimits_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}$ (4.10)
@@
$\sumop\displaylimits_{m=1}^{n-y}\sumop\displaylimits_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}\underbrace{v_{n-m}v_{n-k}}_{\geq(\lambda-\epsilon)^{2}}\geq(\lambda-\epsilon)^{2}\sumop\displaylimits_{m=1}^{n-y}\sumop\displaylimits_{k=1}^{m-y}p_{m}^{*}p_{k}^{*}~.$ (4.11)
@@
We now turn to the proof of Proposition 11, which asserts that the number of leaves in the medium-tail scenario diverges in probability. We have already shown that $J_{n}\to\infty$ in probability, where $J_{n}$ counts the number of edges connecting vertices in the chain $S_{n}$ to leaves via their longest outgoing edge. Proposition 11 follows from this result, but some care is needed: $J_{n}$ counts edges, not distinct leaves, and multiple edges may connect to the same leaf. The key idea of the proof is to show that the number of edges of the type $\max(Z_{n},W_{n})$ connecting to any single leaf cannot grow to infinity. This ensures that the divergence of $J_{n}$ indeed implies that the number of distinct leaves connected to $T_{n}$ also diverges to infinity in probability.
@@
$\leq\mathbb{P}\left(\left[\max_{k\in\mathcal{L}_{n}^{*}}\sumop\displaylimits_{m\in S_{n}}\mathbbm{1}_{\max(Z_{m},W_{m})=m-k}<\frac{\mathbb{E}J_{n}}{2x}\right]\cap\left[\left|\mathcal{L}_{n}^{*}\right|<x\right]\right)$
@@
$J_{n}=\sumop\displaylimits_{k\in\mathcal{L}_{n}^{*}}\sumop\displaylimits_{m\in S_{n}}\mathbbm{1}_{\max(Z_{m},W_{m})=m-k}~,$
@@
$\mathbb{P}\left(\left[\max_{k\in\mathcal{L}_{n}^{*}}\sumop\displaylimits_{m\in S_{n}}\mathbbm{1}_{\max(Z_{m},W_{m})=m-k}<\frac{\mathbb{E}J_{n}}{2x}\right]\cap\left[\left|\mathcal{L}_{n}^{*}\right|<x\right]\right)$
@@
$\mathbb{P}(D_{n,k}>y)$ $=\mathbb{P}\left(\sumop\displaylimits_{m=1}^{n}\mathbbm{1}_{\max(Z_{m},W_{m})=m-k}>y\right)$
@@
$\sumop\displaylimits_{m=1}^{\infty}(p_{m}^{*})^{y_{n}}\leq\sumop\displaylimits_{m=1}^{\infty}(p_{m}^{*})^{2}<\infty\text{ (since }\mathbb{E}\min(Z,W)<\infty\text{ )}~.$
@@
$\mathbb{P}\left(\cup_{k\leq 0}\left[D_{n,k}\geq\frac{\mathbb{E}J_{n}}{2x}\right]\right)$ $\leq\sumop\displaylimits_{k\leq 0}\left(\frac{e}{y_{n}}\right)^{y_{n}}(p_{1-k}^{*})^{y_{n}}$
$\leq\left(\frac{e}{y_{n}}\right)^{y_{n}}\sumop\displaylimits_{k=1}^{\infty}(p_{k}^{*})^{2}\to 0~.$ (4.15)
@@
Having established divergence in probability, we are now ready to prove almost sure divergence.
@@
$\mathbb{P}\left(Y_{n}\leq k\right)\leq\sumop\displaylimits_{m\geq n}\mathbb{P}\left(\min(Z_{m},W_{m})>m-k\right).$
Since $\mathbb{E}\min(Z,W)<\infty$, the sum above is finite and goes to 0 as $n$ goes to infinity. Hence, $Y_{n}\to\infty$ in probability.
Let us fix $n\geq 0$ and apply Corollary 9 to the model obtained by shifting by $n$ the set of integers so that leaves are indexed by the set $\{k:k\leq n\}$ and the edges are given by the sequences of random variables $(Z_{k})_{k\geq n+1}$ and $(W_{k})_{k\geq n+1}$. This gives us a random variable $N_{n}>n$, finite almost surely, such that $N_{n}\in S_{k}$ for all $k\geq N_{n}$. Let $I_{n}\leq n\}$ be the root of this tree, and note that